# Entanglement dynamics in double Jaynes-Cummings model and intensity-dependent double Jaynes-Cummings model for squeezed coherent thermal states

### v3

@@ 8 Effects of dipole-dipole interaction between the two atoms
People have shown interest in studying the effects of the dipole-dipole interaction on the dynamics of the atom-field system for a long time. In particular, the authors in [81], have discussed a detailed dynamics of the two-atom system where two atoms are interacting with each other via dipole-dipole coupling. A remarkable feature which comes out of this study is that more series of revivals appear in the dynamics that are like the one-atom case, partial and overlapping. They have also found that quantum collapse is no longer Gaussian and now depends on the dipole-dipole interaction parameter. In ref.[82], Evseev et. al, have investigated entanglement dynamics between two qubits in a non-resonant DJCM taking into account the direct dipole-dipole interaction between the qubits. The results show that the dipole-dipole parameter has great impact on the entanglement dynamics. They have also shown that the presence of large $g_{d}$ leads to a stabilization of entanglement for all Bell-types initial qubit states and different couplings and detunings. In another paper[83] the authors have studied the nonlinear version of the Jaynes-Cummings model for two identical two-level atoms for Ising-like and dipole-dipole interaction between the atoms. The have found that when the ratio of $\frac{g_{d}-J_{z}}{\lambda}>>1$, some significant results are observed. The effects of dipole-dipole interaction on the energy levels of an one dimensional system using Jaynes-Cummings Hubbard model is discussed in Ref.[84].
@@ 9 Conclusion
In this work we have investigated the entanglement dynamics of the atom-atom, atom-field and field-field subsystems with squeezed coherent thermal states in intensity-dependent double Jaynes-Cummings model (IDDJCM) and double Jaynes-Cummings model (DJCM). It is noticed that the atom-atom and field-field entanglements almost remain unchanged with increasing number of squeezed photons while keeping coherent and thermal photons constant for IDDJCM. An array of entanglement pulses is formed with almost same ESDs and same peak amplitude. However, the atom-field entanglements are affected noticeably. For DJCM, all the dynamics change with increasing squeezed photons and length of ESDs also increases for atom-atom and atom A-field b subsystems.

### v4

@@ 8 Effects of dipole-dipole interaction between the two atoms
Much interest have been shown in studying the effects of the dipole-dipole interaction on the dynamics of the atom-field system for a long time. In particular, the authors in [81], have discussed a detailed dynamics of the two-atom system where two atoms are interacting with each other via dipole-dipole coupling. A remarkable feature which comes out of this study is that more series of revivals appear in the dynamics that are like the one-atom case, partial and overlapping. They have also found that quantum collapse is no longer Gaussian and now depends on the dipole-dipole interaction parameter. In Ref.[82], Evseev et. al., have investigated entanglement dynamics between two qubits in a non-resonant DJCM taking into account the direct dipole-dipole interaction between the qubits. The results show that the dipole-dipole parameter has great impact on the entanglement dynamics. They have also shown that the presence of large $g_{d}$ leads to a stabilization of entanglement for all Bell-types initial qubit states and different couplings and detunings. In another paper[83] , the authors have studied the nonlinear version of the Jaynes-Cummings model for two identical two-level atoms for Ising-like and dipole-dipole interaction between the atoms. The have found that when the ratio of $\frac{g_{d}-J_{z}}{\lambda}>>1$, some significant results are observed. The effects of dipole-dipole interaction on the energy levels of an one dimensional system using Jaynes-Cummings Hubbard model is discussed in Ref.[84].
@@ 9 Conclusion
In this work, we have investigated the entanglement dynamics of the atom-atom, atom-field and field-field subsystems with squeezed coherent thermal states in intensity-dependent double Jaynes-Cummings model (IDDJCM) and standard double Jaynes-Cummings model (DJCM). It is noticed that the atom-atom and field-field entanglements almost remain unchanged with increasing number of squeezed photons while keeping coherent and thermal photons constant for IDDJCM. An array of entanglement pulses is formed with almost same ESDs and same peak amplitude. However, the atom-field entanglements are affected noticeably. For DJCM, all the dynamics change with increasing squeezed photons and length of ESDs also increases for atom-atom and atom A-field b subsystems.